# MiniMaxAD: A Lightweight Autoencoder for Feature-Rich Anomaly Detection

### v3

@@ 4.2 Anomaly Detection and Segmentation on FPADs
GoodsAD (single-class) PatchCore 85.9 - - - - - 85.7 -
RD 66.5 - - - - - 79.7 -
SimpleNet 75.3 - - - - - 68.8 -
MiniMaxAD-fr(Ours) 86.1 - - - - - 87.5 -
4.3 Anomaly Detection and Segmentation on FRADs
We compare our method with current SoTA methods using image-level and pixel-level metrics (see Section 4.1) across a range of datasets. This section focuses on comparisons with UniAD(You et al. 2022), DiAD(He et al. 2024b), and MambaAD(He et al. 2024a), which have been recently tailored for multi-class settings. We also compare other SoTA methods, such as RD(Deng and Li 2022), SimpleNet(Liu et al. 2023) and PatchCore(Ishida et al. 2023). Additionally, a comparison with a more recent MVAD(He et al. 2024c) method designed for multi-view anomaly detection is included.
As shown in Table 3, our MiniMaxAD-fr outperforms all comparative methods on FRADs. Specifically, in the multi-class benchmark, our method’s I-AUROC improves by 0.2$\uparrow$/2.6$\uparrow$ on the MVTecAD and VisA datasets, respectively, compared to MambaAD. On the mAD metric, we see an improvement of 0.1$\uparrow$/3.6$\uparrow$. When evaluated on the more challenging Real-IAD dataset, our mAD metrics are on par with MambaAD, both reaching 72.7. In the single-class Real-IAD benchmark, our method improves detection performance at the image level by 1.0$\uparrow$/1.5$\uparrow$/1.4$\uparrow$ and the mAD metrics by 5.0$\uparrow$ compared to MVAD, which is specifically designed for multi-view settings. Additionally, the single-class GoodsAD further demonstrates the competitiveness of our method, achieving a 0.2$\uparrow$/1.8$\uparrow$ improvement in the two fundamental metrics, I-AUROC and AUPRO, compared to PatchCore. Furthermore, our method shows a significant efficiency advantage. These results fully demonstrate the validity and necessity of our approach. More detailed results are presented in the Appendix.
FRADs FPADs
GoodsAD Uni ViSA Uni AeBAD-S&MVTecAD AeBAD-S ViSA
RD RD + $\mathcal{L}_{global}$ 66.5 92.7 76.2&89.6 81.0 96.0
Ours + $\mathcal{L}_{local}$ 80.1 93.0 84.3&97.1 85.1 93.8
Ours + $\mathcal{L}_{local-hm}$ 80.6 96.3 86.2&98.2 85.4 96.0
Ours-fp Ours + $\mathcal{L}_{global}$ 77.5 93.1 84.9&96.3 87.4 96.1
Ours-fr Ours + $\mathcal{L}_{adc}$ 86.1 96.9 87.2&98.4 83.9 95.6
@@ 4.3 Anomaly Detection and Segmentation on FRADs
Table 5: Ablation study on structural elements of the model.
Method Params(M) FLOPs(G) FPS mAD
RD 117.1 30.7 665 76.2
UniAD 24.5  3.6 188 74.6
SimpleNet 72.8 16.1 52 65.3
DiAD 1331.3 451.5 - 68.9
MambaAD 25.7 8.3 216 79.1
Ours 33.1 7.0 906 80.4
@@
Table 6: Complexity comparison of SoTA methods. Frames Per Second (FPS) are measured on an NVIDIA RTX 4090 with a batch size of 16. mAD represents the mean value of mAD across the benchmark for the three multi-class settings.
@@ 4.4 Ablation and Analysis
Comparison of the effectiveness of the model’s structural elements. Table 5 presents ablation studies on the Uni MVTecAD dataset for GRN and large kernel designs. Specifically, we pre-trained four encoders on a sampled version of ImageNet, ImageNet-100(Tian, Krishnan, and Isola 2019), and their validation set classification scores were nearly identical. Further details on the pre-training process are provided in Appendix. For the same architecture, models using GRN typically outperform those without GRN. Additionally, performance tends to decline and lacks significant comparability when the SarK Block is used as an encoder.
@@
Comparison of loss functions applied to different types of datasets. A comprehensive ablation study is presented in Table 4, validating the effectiveness of the loss function and dataset taxonomy. Compared to RD, MiniMaxAD+$\mathcal{L}_{global}$ demonstrates a significant improvement in performance on FRADs and FPADs, affirming the suitability of large kernel convolutional networks with GRN units for the multi-scale reconstruction framework. Our asymptotic adaptive contraction strategy, which is tailored for FRADs based on $\mathcal{L}_{local}$, achieves optimal performance on FRADs. In contrast, the performance of MiniMaxAD+$\mathcal{L}_{adc}$ is degraded compared to MiniMaxAD+$\mathcal{L}_{global}$ on FPADs. The reason for this degradation is that while $\mathcal{L}_{adc}$ is based on $\mathcal{L}_{local}$ and offers certain advantages, it compromises global consistency to some extent(Guo et al. 2023). Specifically, easier-to-fit feature-poor anomaly detection datasets (FPADs) do not benefit from $\mathcal{L}_{adc}$, which relies on point-by-point distances, as these distances are more stringent compared to global distances. However, transforming FPADs into FRADs is straightforward. As reported in the third column of our table, we merged scenarios to unify AeBAD-S with MVTec AD for training. Here we resize MVTec AD to 224 $\times$ 224, other settings follow the default AeBAD-S. The results demonstrate that converting FPADs to FRADs using the unified setting in our methodology yields more gains than losses. The additional results in Table 1 further highlight the significant improvement in segmentation performance. Further ablation studies, experimental results, and qualitative visualizations are presented in the Appendix.
@@
5 Conclusion
In this study, we revisit existing UAD datasets and categorize them into feature-rich anomaly detection datasets (FRADs) and feature-poor anomaly detection datasets (FPADs). We analyze the significant performance degradation of previous state-of-the-art (SoTA) methods when processing FRADs from the perspective of information capacity. To address this, we introduce MiniMaxAD, which enhances the “effective capacity” of the network by focusing on two key elements: feature diversity enhancement and large kernel convolution. These improvements allow for the efficient modeling of large amounts of information using smaller models. In addition, we introduce the hard mining loss ADCLoss, featuring a highly adaptive discard rate, to further enhance the model’s ability to handle FRADs. Extensive testing on a range of mature and challenging UAD datasets confirms the superiority of our approach, demonstrating its potential to advance the development of multiscale reconstruction-based methods.
Limitation. In this work, we avoided incorporating additional optimizations, such as the simulation of anomalies or the migration of pre-trained encoders from specific domains, in order to maintain efficiency and scalability. Additionally, MiniMaxAD’s smaller boost in processing FPADs compared to RDs may be a limitation, but in our methodology, converting FPADs to FRADs using the unified setting is a superior option, which improves performance while significantly reducing storage.
Appendix A Datasets
@@
VisA is a challenging anomaly detection dataset with 10,821 images across 12 subsets, each representing a different class of objects. The anomalies in these images include surface defects like scratches, dents, discoloration, or cracks, and structural defects such as misaligned or missing parts. We applied the unsupervised default settings (Zou et al. 2022) to segregate the training and test sets. In our experiments, all images were resized to a uniform resolution of 256 $\times$ 256.
@@
Table 7: Score of the pre-trained encoder on the validation set.
Appendix B Additional Details
Data sources. The results for the comparison methods on AeBAD-S (Zhang et al. 2023b) and GoodsAD (Zhang et al. 2024) are partially from the original paper. The results of the MVTecAD, ViSA, and Real-IAD comparison tests in the multi-class setting were extracted from MambaAD (He et al. 2024a). Meanwhile, the comparison data for Real-IAD in the single-class setting was partially taken from MVAD (He et al. 2024c). We are very grateful to the authors for conducting these detailed experiments.
Pre-trained encoder. We pre-trained four encoders on a subset of ImageNet, specifically ImageNet-100(Tian, Krishnan, and Isola 2019). For the 100 categories in the training set, we randomly selected 100 images per category to form the validation set. Each of our models was trained for 50 epochs, and their validation set scores are presented in Table A1. At the model level, the LarK block refers to the original UniRepLKNet-N architecture, where the large kernel block is the DilatedReparamBlock with a kernel size of 13. The SarK module, on the other hand, replaces the DilatedReparamBlock with a 3x3 depth-wise convolution.
Appendix C Additional Experiments and Analysis
@@ C.1 Alignment Process in Entropy Perspective
$H(\mathcal{F})=-\sum_{i=1}^{n}p(\mathcal{f}_{i})\log_{2}(p(\mathcal{f}_{i})+\epsilon)$
@@ C.2 Additional Stability Experiments
The results in the main paper are reported with a single random seed following our baseline (Deng and Li 2022). In Tables 8 and 9, we report the mean and standard deviation of three runs with three different random seeds (1, 11 and 111) (Guo et al. 2023).
@@
Table 8: Performance on AeBAD-S over three runs.
@@
Table 9: Performance on GoodsAD over three runs.
@@ C.3 Additional Ablation Study
Appendix D More Quantitative Results
More quantitative results on the GoodsAD dataset. Table 12 presents the image-level anomaly detection results and pixel-level anomaly localization quantitative results for all categories in the GoodsAD dataset under the separated single-class setting
@@ Appendix D More Quantitative Results
Appendix E Qualitative Results
In Figure 6, we display anomaly maps for 14 samples from GoodsAD, and in Figure 7, we present anomaly maps for 11 samples each from Uni MVTec AD and Uni ViSA. To ensure that the visualization results effectively highlight the key qualities of the methods, we employed Min-Max normalization for all anomaly maps produced by the different methods.
RD SimpleNet PatchCore Our-fr
cigarette box 65.1/72.7 97.8/90.3 98.0/92.7 98.1/86.3
drink bottle 67.2/84.2 70.0/66.5 83.5/85.9 80.4/86.7
drink can 63.3/81.9 70.8/67.0 93.4/89.6 90.3/87.6
food bottle 76.6/84.3 78.1/75.0 89.1/88.9 91.8/92.8
food box 63.3/75.9 72.0/56.4 74.5/75.4 77.8/84.1
food package 63.5/79.4 63.2/57.8 77.0/81.6 77.9/87.2
Average 66.5/79.7 75.3/68.8 85.9/85.7 86.1/87.5
Table 12: Comparison with SoTA methods on MVTec-AD dataset for single-class anomaly detection and segmentation with I-AUROC/AUROC metrics.
@@ Appendix E Qualitative Results
Bottle 99.6 /99.9 /98.4 99.7 /100./100. 100. /100./100. 99.7 /96.5 /91.8 100. /100. 100. /100. /100.
Cable 84.1 /89.5 /82.5 95.2 /95.9 /88.0 97.5 /98.5 /94.7 94.8 /98.8 /95.2 98.8/99.2 99.2 /99.6 /96.3
Capsule 94.1 /96.9 /96.9 86.9 /97.8 /94.4 90.7 /97.9 /93.5 89.0 /97.5 /95.5 94.4/98.7 96.7 /99.3 /96.3
Hazelnut 60.8 /69.8 /86.4 99.8 /100./99.3 99.9 /99.9 /99.3 99.5 /99.7 /97.3 100. /100. 100. /100. /100.
MetalNut 100. /100. /99.5 99.2 /99.9 /99.5 96.9 /99.3 /96.1 99.1 /96.0 /91.6 99.9 /100. 100. /100. /100.
Pill 97.5/99.6 /96.8 93.7 /98.7 /95.7 88.2 /97.7 /92.5 95.7 /98.5 /94.5 97.0 /99.5 97.7 /99.6 /97.9
Screw 97.7 /99.3/95.8 87.5 /96.5 /89.0 76.7 /90.6 /87.7 90.7 /99.7 /97.9 94.7 /97.9 95.6/98.6 /94.0
Toothbrush 97.2 /99.0 /94.7 94.2 /97.4 /95.2 89.7 /95.7 /92.3 99.7 /99.9 /99.2 98.3 /99.3 98.6/99.5/96.8
Transistor 94.2 /95.2 /90.0 99.8/98.0 /93.8 99.2 /98.7 /97.6 99.8/99.6/97.4 100. /100. 98.2 /97.7 /91.1
Zipper 99.5 /99.9 /99.2 95.8 /99.5 /97.1 99.0 /99.7 /98.3 95.1 /99.1 /94.4 99.3 /99.8 99.5 /99.9 /98.7
Carpet 98.5 /99.6 /97.2 99.8 /99.9 /99.4 95.7 /98.7 /93.2 99.4 /99.9 /98.3 99.8 /99.9 99.6 /99.9 /98.9
Grid 98.0 /99.4 /96.5 98.2 /99.5 /97.3 97.6 /99.2 /96.4 98.5/99.8/97.7 100. /100. 98.1 /99.5 /98.2
Leather 100. /100. /100. 100. /100./100. 100. /100./100. 99.8 /99.7 /97.6 100. /100. 100. /100. /100.
Tile 98.3 /99.3 /96.4 99.3/99.8 /98.2 99.3/99.8 /98.8 96.8 /99.9/98.4 98.2 /99.3 100. /100. /100.
Wood 99.2/99.8/98.3 98.6 /99.6 /96.6 98.4 /99.5 /96.7 99.7 /100. /100. 98.8 /99.6 99.2/99.8/97.5
Mean 94.6 /96.5 /95.2 96.5 /98.8 /96.2 95.3 /98.4 /95.8 97.2 /99.0 /96.5 98.6/99.6 98.8 /99.6 /97.7
@@
Figure 6: Anomaly maps on anomalous images from GoodsAD.
Figure 7: Anomaly maps on anomalous images from MVTec AD and ViSA.

### v4

@@ 4 Experiments
GoodsAD (single-class) PatchCore 85.2 85.9 82.9 97.2 49.4 53.0 84.9 76.9
RD 67.7 68.9 74.1 93.3 16.4 23.8 79.0 60.5
SimpleNet 74.1 76.9 75.5 86.9 21.9 27.8 65.2 61.2
MiniMaxAD-fr(Ours) 86.2 86.9 82.7 96.2 43.9 46.9 86.0 75.5
Figure 5: Visualization of qualitative results for the MVTecAD dataset in the multi-class setting.
4.1 Experimental Settings
Datasets. MVTec AD and VisA feature high-resolution images across multiple object and texture categories, highlighting both surface and structural defects. They are the two most popular datasets for anomaly detection, covering a wide range of anomalies at different scales in scenarios such as components, food, and medical care. AeBAD-S is specifically focused on domain diversity within aero-engine blade images.The GoodsAD dataset addresses product damage detection in unmanned supermarket scenarios. It captures extensive intra-class diversity with 6,124 high-resolution images of common supermarket goods, showcasing a single product offered by multiple brands. Real-IAD is a recent, large industrial anomaly detection dataset comprising 150K images across 30 categories, with each category containing five default perspectives. Detailed information is presented in Appendix.
Implementation. Network optimization was carried out using an AdamW optimizer [23], with the $\beta$ parameters set to (0.9, 0.999) and a weight decay of 0.00005. The learning rate was set at 0.005, accompanied by an ExponentialLR scheduler with a gamma of 0.995. The batch size for all experiments was 16. For experiments in the single-class setting, GoodsAD was trained for 240 epochs, ViSA for 150 epochs, AeBAD-S for 130 epochs, and Real-IAD for 50 epochs. In the multi-class setting, MVTec AD and ViSA were trained for 160 epochs, while Real-IAD was trained for only 1 epoch. Results were reported from the last epoch for each scenario. All experiments were performed on an NVIDIA RTX 4090 GPU using PyTorch 2.0. The model does not use efficient large kernel convolution as suggested by [10].
Metrics. We use two of the most commonly employed metrics for anomaly detection and segmentation as our basic evaluation metrics [2]: Area Under the Receiver Operating Characteristic Curve (AUROC) and pixel-level Area Under the Per-Region Overlap (AUPRO). Additionally, following previous work [40, 15], we adopt supplementary metrics on certain datasets for a more comprehensive evaluation, including Average Precision (AP) [39] and F1-score-max (F1_max) [44]. We also calculate the average of the seven evaluation metrics mentioned above, denoted as mAD, to represent the model’s overall performance.
4.2 Anomaly Detection and Segmentation on FPADs
Table 4: Ablation study on structural elements of the model.
@@ 4.2 Anomaly Detection and Segmentation on FPADs
In this section, we focus on demonstrating the performance of MiniMaxAD on FPADs using the two most commonly used metrics (see Section 4.1). We compare our method with RD [8], PatchCore [18], MMR [43], EfficientAD [1], and MambaAD [15], all of which are powerful state-of-the-art (SoTA) methods.
It is important to emphasize that anomaly detection for FPADs is not the primary focus of our research. While some methods, such as MambaAD, perform exceptionally well on FRADs, they may still struggle with FPADs. This limitation precisely motivates our proposed classification approach, which aims to identify targeted solutions based on different application scenarios.
Figure 6: Visualization of qualitative results for the GoodsAD dataset in the single-class setting.
FRADs FPADs
GoodsAD Uni ViSA Uni AeBAD-S&MVTecAD AeBAD-S ViSA
Baseline RD + $\mathcal{L}_{global}$ 66.5 92.7 76.2&89.6 81.0 96.0
Ours + $\mathcal{L}_{local}$ 80.1 93.0 84.3&97.1 85.1 93.8
Ours + $\mathcal{L}_{local-hm}$ 80.6 96.3 86.2&98.2 85.4 96.0
Ours-fp Ours + $\mathcal{L}_{global}$ 77.5 93.1 84.9&96.3 87.4 96.1
Ours-fr Ours + $\mathcal{L}_{adc}$ 86.1 96.9 87.2&98.4 83.9 95.6
4.3 Anomaly Detection and Segmentation on FRADs
Table 6: Ablation study on on the pre-trained backbone. For the depth column, “9+9” means 9 LarK Blocks and 9 SmaK Blocks. Our model uses UniRepLKNet-N by default.
Backbone Depth Params(M) Image-level Pixel-level mAD
AUROC AP F1_max AUROC AP F1_max AUPRO
UniRepLKNet-F $\left[2,2,6+0\right]$ 10.8 97.3 98.7 96.7 94.7 51.3 54.0 89.5 83.2
UniRepLKNet-P $\left[2,2,6+0\right]$ 19.0 98.1 99.3 97.5 95.0 53.4 56.0 90.7 84.3
UniRepLKNet-N $\left[2,2,8+0\right]$ 33.1 98.8 99.6 97.7 96.1 58.5 59.5 92.2 86.1
UniRepLKNet-T $\left[2,2,9+9\right]$ 51.4 99.0 99.6 98.0 97.1 61.5 61.9 93.6 87.2
UniRepLKNet-S $\left[2,2,9+18\right]$ 96.4 99.1 99.7 98.3 97.1 62.8 62.7 93.8 87.6
@@ 4.3 Anomaly Detection and Segmentation on FRADs
Table 7: Complexity comparison of SoTA methods. All complexity metrics are measured on an NVIDIA RTX 4090 with a batch size of 16 , and the input resolution is $256\times 256$. The CPU used is an AMD EPYC 9654 96-Core Processor with 16 vCPUs allocated. mAD represents the mean value of mAD across the benchmark for the three multi-class settings.
Method Params(M) FLOPs(G) FPS Memory(M) Storage(M) mAD
RD 117.1 30.7 665 2179 448 76.2
UniAD 24.5  3.6 188 1373 95 74.6
SimpleNet 72.8 16.1 52 1035 279 65.3
DiAD 1331.3 451.5 - - - 68.9
MambaAD 25.7 8.3 216 2309 99 79.1
Ours 33.1 7.0 906 1327 127 80.4
Quantitative results. We compare our method with current SoTA methods using image-level and pixel-level metrics (see Section 4.1) across a range of datasets. This section focuses on comparisons with UniAD [37], DiAD [14], and MambaAD [15], which have been recently tailored for multi-class settings. We also compare other SoTA methods, such as RD [8], SimpleNet [22] and PatchCore [18]. Additionally, a comparison with a more recent MVAD [16] method designed for multi-view anomaly detection is included.
As shown in Table 3, our MiniMaxAD-fr outperforms all comparative methods on FRADs. Specifically, in the multi-class benchmark, our method’s I-AUROC improves by 0.2$\uparrow$/2.6$\uparrow$ on the MVTecAD and VisA datasets, respectively, compared to MambaAD. On the mAD metric, we see an improvement of 0.1$\uparrow$/3.6$\uparrow$. When evaluated on the more challenging Real-IAD dataset, our mAD metrics are on par with MambaAD, both reaching 72.7. In the single-class Real-IAD benchmark, our method improves detection performance at the image level by 1.0$\uparrow$/1.5$\uparrow$/1.4$\uparrow$ and the mAD metrics by 5.0$\uparrow$ compared to MVAD, which is specifically designed for multi-view settings. Additionally, the single-class GoodsAD further demonstrates the competitiveness of our method, achieving a 1.0$\uparrow$/1.1$\uparrow$ improvement in the two fundamental metrics, I-AUROC and AUPRO, compared to PatchCore. While the mAD of MiniMaxAD-fr in GoodsAD is lower than that of PatchCore, which is challenging to implement in multi-class setting due to its memory bank mechanism, our approach remains scalable. These results fully demonstrate the validity and necessity of our approach. More detailed results are presented in the Appendix.
Qualitative results. We visualize the experimental results of MiniMaxAD-fr compared to the baseline model RD in the multi-class setting to further illustrate the model’s lack of “effective capacity”, as shown in Figure 5. Specifically, during the training process of the reconstruction model, the objective is to iteratively train the model to reconstruct normal samples from the training set. In the multi-class setting, sharing a single model across numerous items significantly exacerbates the model’s memory burden. Although the RD model has more than three times the number of parameters as MiniMaxAD, it struggles to accurately memorize the normal regions of objects, resulting in numerous false-negative regions in the anomaly map. More concerning, its inconsistent memory even causes the anomaly maps to display unexpected outliers, as seen in the bottom regions of the first four columns of the anomaly maps for the MVTecAD dataset in Figure 5.
Figure 6 presents the qualitative results on GoodsAD, along with a comparison to PatchCore. As shown in the figure, PatchCore benefits from its explicit memory bank memorization mechanism, which prevents misremembering and results in smoother anomaly maps overall. However, this explicit memorization becomes a significant burden as the training set increases in size. Additionally, The figure shows that PatchCore encounters some false positives when dealing with domain shift, and the performance degradation presented in Table 1 directly reflects this issue.
@@ 4.4 Ablation and Analysis
Comparison of the effectiveness of the model’s structural elements. Table 4.2 presents ablation studies on the Uni MVTecAD dataset for GRN and large kernel designs. Specifically, we pre-trained four encoders on a sampled version of ImageNet, ImageNet-100 [29], and their validation set classification scores were nearly identical. Further details on the pre-training process are provided in Appendix. For the same architecture, models using GRN typically outperform those without GRN. Additionally, performance tends to decline and lacks significant comparability when the SarK Block is used as an encoder.
@@
Comparison of loss functions applied to different types of datasets. A comprehensive ablation study is presented in Table 5, validating the effectiveness of the loss function and dataset taxonomy. Compared to RD, MiniMaxAD+$\mathcal{L}_{global}$ demonstrates a significant improvement in performance on FRADs and FPADs, affirming the suitability of large kernel convolutional networks with GRN units for the multi-scale reconstruction framework. Our asymptotic adaptive contraction strategy, which is tailored for FRADs based on $\mathcal{L}_{local}$, achieves optimal performance on FRADs. In contrast, the performance of MiniMaxAD+$\mathcal{L}_{adc}$ is degraded compared to MiniMaxAD+$\mathcal{L}_{global}$ on FPADs. The reason for this degradation is that while $\mathcal{L}_{adc}$ is based on $\mathcal{L}_{local}$ and offers certain advantages, it compromises global consistency to some extent [13]. Specifically, easier-to-fit feature-poor anomaly detection datasets (FPADs) do not benefit from $\mathcal{L}_{adc}$, which relies on point-by-point distances, as these distances are more stringent compared to global distances. However, transforming FPADs into FRADs is straightforward. As reported in the third column of our table, we merged scenarios to unify AeBAD-S with MVTec AD for training. Here we resize MVTec AD to 224 $\times$ 224, other settings follow the default AeBAD-S. The results demonstrate that converting FPADs to FRADs using the unified setting in our methodology yields more gains than losses. The additional results in Table 1 further highlight the significant improvement in segmentation performance. Further ablation studies and experimental results are presented in the Appendix.
Figure 7: Comparison of the ERF of different networks.
Comparison of different pre-trained backbones & receptive field analysis. According to the Effective Receptive Field (ERF) theory [24], a model’s effective receptive field occupies only a fraction of its theoretical receptive field and exhibits Gaussian distribution characteristics. The extent of this influence can be determined by calculating the gradient of a point on the deep feature map with respect to the original pixel. Recent studies [19, 9] have visualized ERFs by backpropagating from the center of the deep feature map. Following the visualization approach described in [9], Figure 7 presents the ERF visualization results for various backbones [10, 38]. As shown in the figure, deeper models exhibit larger ERFs, indicating that the bottleneck layer can integrate more abstract features. Additionally, we evaluated the impact of backbones with larger ERFs on model performance using the Uni MVTecAD dataset. Across all backbones, our bottleneck layer consistently employs two Smark modules, and the decoder mirrors the encoder directly. The results in Table 6 demonstrate that our model adheres to the “scaling law”, suggesting that larger backbones enhance performance. Notably, even when using the larger UniRepLKNet-S as the backbone, the total number of model parameters remains lower than that of the standard RD.
@@
5 Conclusions
This paper introduces SSFilter, which is the first method to apply sample-level filtering in noisy unsupervised anomaly detection. The advantage of this filtering approach is its high scalability. SSFilter not only achieves impressive performance in end-to-end fully unsupervised anomaly detection but also allows filtering of the entire dataset using a trained model, thus serving as a bridge between UAD and FUAD methods. We conducted extensive experiments to demonstrate the outstanding performance of our approach.
limitations. Despite SSFilter’s strong performance in noisy scenarios, its filtering mechanism discards some normal samples, causing a degree of performance degradation. Data-efficient unsupervised anomaly detection methods could improve our approach in future work.
Supplementary Material
A Datasets
@@
VisA is a challenging anomaly detection dataset with 10,821 images across 12 subsets, each representing a different class of objects. The anomalies in these images include surface defects like scratches, dents, discoloration, or cracks, and structural defects such as misaligned or missing parts. We applied the unsupervised default settings [44] to segregate the training and test sets. In our experiments, all images were resized to a uniform resolution of 256 $\times$ 256.
@@
B Additional Details
Data sources. The results for the comparison methods on AeBAD-S [43] are partially from the original paper. The results of the MVTecAD, ViSA, and Real-IAD comparison tests in the multi-class setting were extracted from MambaAD [15]. Meanwhile, the comparison data for Real-IAD in the single-class setting was partially taken from MVAD [16]. We are very grateful to the authors for conducting these detailed experiments.
Pre-trained encoder. We pre-trained four encoders on a subset of ImageNet, specifically ImageNet-100 [29]. For the 100 categories in the training set, we randomly selected 100 images per category to form the validation set. Each of our models was trained for 50 epochs, and their validation set scores are presented in Table A1. At the model level, the LarK block refers to the original UniRepLKNet-N architecture, where the large kernel block is the DilatedReparamBlock with a kernel size of 13. The SarK module, on the other hand, replaces the DilatedReparamBlock with a 3x3 depth-wise convolution.
C Additional Experiments and Analysis
@@ C.1 Alignment Process in Entropy Perspective
$H(\mathcal{F})=-\sum_{i=1}^{n}p(\mathcal{f}_{i})\log_{2}(p(\mathcal{f}_{i})+\epsilon)$ (10)
@@ C.2 Additional Stability Experiments
The results in the main paper are reported with a single random seed following our baseline [8]. In Tables 9 and 10, we report the mean and standard deviation of three runs with three different random seeds (1, 11 and 111) [13].
Table 9: Performance on AeBAD-S over three runs.
@@
Table 10: Performance on GoodsAD over three runs.
@@ C.3 Additional Ablation Study
D More Quantitative Results
@@
Bottle 99.6 /99.9 /98.4 99.7 /100./100. 100. /100./100. 99.7 /96.5 /91.8 100. /100./100. 100. /100. /100.
Cable 84.1 /89.5 /82.5 95.2 /95.9 /88.0 97.5 /98.5 /94.7 94.8 /98.8 /95.2 98.8/99.2/95.7 99.2 /99.6 /96.3
Capsule 94.1 /96.9 /96.9 86.9 /97.8 /94.4 90.7 /97.9 /93.5 89.0 /97.5 /95.5 94.4/98.7/94.9 96.7 /99.3 /96.3
Hazelnut 60.8 /69.8 /86.4 99.8 /100./99.3 99.9 /99.9 /99.3 99.5 /99.7 /97.3 100. /100./100. 100. /100. /100.
MetalNut 100. /100. /99.5 99.2 /99.9 /99.5 96.9 /99.3 /96.1 99.1 /96.0 /91.6 99.9 /100./99.5 100. /100. /100.
Pill 97.5/99.6 /96.8 93.7 /98.7 /95.7 88.2 /97.7 /92.5 95.7 /98.5 /94.5 97.0 /99.5/96.2 97.7 /99.6 /97.9
Screw 97.7 /99.3/95.8 87.5 /96.5 /89.0 76.7 /90.6 /87.7 90.7 /99.7 /97.9 94.7 /97.9/94.0 95.6/98.6 /94.0
Toothbrush 97.2 /99.0 /94.7 94.2 /97.4 /95.2 89.7 /95.7 /92.3 99.7 /99.9 /99.2 98.3 /99.3/98.4 98.6/99.5/96.8
Transistor 94.2 /95.2 /90.0 99.8/98.0 /93.8 99.2 /98.7 /97.6 99.8/99.6/97.4 100. /100./100. 98.2 /97.7 /91.1
Zipper 99.5 /99.9 /99.2 95.8 /99.5 /97.1 99.0 /99.7 /98.3 95.1 /99.1 /94.4 99.3 /99.8/97.5 99.5 /99.9 /98.7
Carpet 98.5 /99.6 /97.2 99.8 /99.9 /99.4 95.7 /98.7 /93.2 99.4 /99.9 /98.3 99.8 /99.9/99.4 99.6 /99.9 /98.9
Grid 98.0 /99.4 /96.5 98.2 /99.5 /97.3 97.6 /99.2 /96.4 98.5/99.8/97.7 100. /100./100. 98.1 /99.5 /98.2
Leather 100. /100. /100. 100. /100./100. 100. /100./100. 99.8 /99.7 /97.6 100. /100./100. 100. /100. /100.
Tile 98.3 /99.3 /96.4 99.3/99.8 /98.2 99.3/99.8 /98.8 96.8 /99.9/98.4 98.2 /99.3/95.4 100. /100. /100.
Wood 99.2/99.8/98.3 98.6 /99.6 /96.6 98.4 /99.5 /96.7 99.7 /100. /100. 98.8 /99.6/96.6 99.2/99.8/97.5
Mean 94.6 /96.5 /95.2 96.5 /98.8 /96.2 95.3 /98.4 /95.8 97.2 /99.0 /96.5 98.6/99.6/97.8 98.8 /99.6 /97.7